A single conservative force F = (5.0x - 11)i N, where x is in meters, acts on a particle moving along an x axis. The potential energy U associated with this force is assigned a value of 25 J at x = 0. (a) What is the maximum positive potential energy? At what (b) negative value and (c) positive value of x is the potential energy equal to zero?
Answer #1

a) F = 5x -11 and F=0 at x=11/5 = 2.2

U = - integral Fdx = -2.5x^2 +11x + 25 because potential is 25J at x=0.

maximum potential will be when F =0,

maximum pE = -2.5*2.2^2 +11*2.2 +25

= 37.1 J answe

b and c) the roots of quadratic equation are -1.65 m and 6.05 m

the negative value is -1.65 m and and positive is 6.05 m